For the year Y, for five countries, one has the following information: GDP (X) = 850 000 currency X / GDP (Y) = 300 000 currency Y / GDP (Z) = 1 500 000 currency Z / GDP (T) = 725 000 currency T / GDP (W) = 1100000 currency W. Basket of products with identical products over the five economies - Basket of Products = 1.25*A + 3.5*B+2*C. Country X (prices in currency X) (Products: A / B / C): 14 / 19 / 21. Country Y (prices in currency Y) (Products: A / B / C) 25 / 32 / 39. Country Z (prices in currency Z) (Products: A / B / C) 8 / 13 / 19. Country T (prices in currency T) (Products: A / B / C) 11 / 16/ 23. Country W (prices in currency W) (Products: A / B / C) 17 / 19 / 26. The correct answer is:a.The value of GDP of country Y in currency X at PPP is 220847 and represents 36.10% of GDP of country X


The value of GDP of country Y in currency X at PPP is ?????? and represents ????%


1
Expert's answer
2022-01-27T13:09:51-0500

"1.25A+3.5B+2C=GDP" 1.25

"X:1.25(14)+3.5(19)+2(21)=126"

"Y:1.25(25)+3.5(32)+2(39)=221.25"


"126X=850,000"

"X=6,746.03"


"221.25Y= 300,000"

"Y=1,355.93"


To get the GDP of Y in currency X at PPP;

"221.25*6,746.03=1,492,559.14"

the GDP 1,492,559.14 represents 76% of GDP of country X.


"(1492559.14-850000)\/850000 *100=76%"